# Differentially Private Synthetic Data via APIs 3: Using Simulators Instead of Foundation Models

### v1

@@ 4.3 Sim-PE with Simulator-generated Data
In this section, we evaluate Sim-PE using a generated dataset from a computer graphics-based renderer on the CelebA dataset. The results, presented in Tables 1 and 3, highlight the following key takeaways:
Sim-PE effectively selects samples that better match the correct classes. Without any information from the private data, the simulator naturally generates images with incorrect class labels (Fig. 3). Consequently, a downstream gender classifier trained on simulator-generated data can at best achieve a trivial accuracy 61.4%—on the test set, the majority class (female) constitutes 61.4%. Building upon this noisy data, Sim-PE iteratively refines the sample selection. Ultimately, Sim-PE selects samples that better align with the target classes (Fig. 3), leading to an accuracy improvement of up to 21.1% (Table 1(a)).
Sim-PE maintains the strong data quality of PE. As shown in Table 1(b), Sim-PE and PE achieve similar FID scores. Unlike in the MNIST experiments (§ 4.2), where Sim-PE significantly improved over PE, the lack of substantial improvement on CelebA can be attributed to two factors. First, on CelebA, PE with foundation models already ranks 3rd among all methods in terms of FID, leaving little room for further gains. Second, in this experiment, Sim-PE is only provided with a fixed dataset generated from the simulator. As seen in Fig. 3, the simulator-generated images exhibit noticeable differences from real CelebA images, such as faces appearing larger. Since Sim-PE in this setting can only select images without modifying them, it cannot correct such discrepancies. Having access to simulator code, as in § 4.2, could potentially alleviate this issue, as Sim-PE could learn to modify the parameters that control the face size. Another potential direction for improvement is a hybrid approach that enables PE to leverage both foundation models and simulators, which we explore preliminarily in the next section.
@@
(c) Sim-PE-generated images ($\epsilon=10$)
Figure 4: The real and generated images on CelebA. The simulator is a weak rule-based avatar generator (Escartín, 2021) significantly different from the real dataset. The top rows correspond to the “female” class, and the bottom rows correspond to the “male” class. The simulator generates images with incorrect classes. Sim-PE tends to generate faces with long hair for the female class and short hair for the male class (correctly), but the generated images have mode collapse issues.
@@ 4.4 Sim-PE with both Simulators and Foundation Models
In this section, we first examine how Sim-PE performs with weak simulators. We again use the CelebA dataset as the private data, but this time, we switch to a rule-based cartoon avatar generator (Escartín, 2021) as the simulator. As shown in Fig. 4, the avatars generated by the simulator differ significantly from the real CelebA images.
Sim-PE with weak simulators still learns useful features. From Table 3, we observe that downstream classifiers trained on Sim-PE with weak simulators achieve poor classification accuracy. However, two interesting results emerge: (1) Despite the significant difference between avatars and real face images, Sim-PE still captures certain characteristics of the two classes correctly. Specifically, Sim-PE tends to generate faces with long hair for the female class and short hair for the male class (Fig. 4). (2) Although the FID scores of Sim-PE are quite poor (Table 3), they still outperform many baselines (Table 1(b)). This can be explained by the fact that, as shown in Gong et al. (2025), when DP noise is high, the training of many baseline methods becomes unstable. This results in images with noisy patterns, non-face images, or significant mode collapse, particularly for DP-NTK, DP-Kernel, and GS-WGAN. In contrast, Sim-PE is training-free, and thus it avoids these issues.
Next, we explore the feasibility of using PE with both foundation models and the weak avatar simulator (§ 3.4). The results are shown in Table 3.
@@
Unlike other state-of-the-art methods that are tied to a specific data synthesizer, this result suggests that PE is a promising framework that can easily combine the strengths of multiple types of data synthesizers.
@@ 4.5 Validating the Design of Sim-PE
How does Sim-PE with simulator-generated data compare to other data selection algorithms? In § 3.3, we discussed two simple alternative solutions for simulator data selection. The comparison is shown in Table 4. As we can see, Sim-PE with iterative data selection outperforms the baselines on most metrics, validating the intuition outlined in § 3.3. However, the clustering approach used in the second baseline still has merit, as it results in a better FID for $\epsilon=10$. This idea is orthogonal to the design of Sim-PE and could potentially be combined for further improvement. We leave this exploration to future work.
@@ 5 Limitations and Future Work
In this paper, we demonstrate the potential of the PE framework for utilizing powerful simulators in DP image synthesis. We believe that the exploration in this paper only scratches the surface of the full potential of this idea. Further extensions include:
@@
    The approach in § 3.3 can be extended beyond simulator-generated datasets, such as public web data. This could potentially further enhance PE’s performance and enable its application in other areas, such as pre-training data selection for private fine-tuning (Yu et al., 2023; Li et al., 2024).
@@
    In this paper, we explore the potential of Sim-PE for images. However, in domains like networking and systems, foundation models are rarer and simulators are more prevalent. Sim-PE could offer even greater potential in these domains.
@@
    The results of Sim-PE for image synthesis are still outperformed by the best baseline. It would be valuable to push the limits of the PE framework further, such as exploring more effective ways to leverage both simulators and foundation models together.
Appendix A Private Evolution
Alg. 1 presents the Private Evolution (PE) algorithm, reproduced from Lin et al. (2023). This algorithm represents the conditional version of PE, where each generated image is associated with a class label. It can be interpreted as running the unconditional version of PE separately for each class.
@@ 1 $S_{\mathrm{syn}}\leftarrow\emptyset$
4       $S_{1}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$
@@
             $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples,S_{t},\sigma,H\right)$
@@
6             $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
@@ 4       $histogram[i]\leftarrow histogram[i]+1$
Appendix B Experimental Details

### v2

@@ 4.3 Sim-PE with Simulator-generated Data
We evaluate Sim-PE using a generated dataset from a computer graphics-based renderer on the CelebA dataset. The results, presented in Tabs. 1 and 3, highlight the following key takeaways:
Sim-PE selects samples that better match target classes. Without private data, the simulator naturally generates images with incorrect labels (Fig. 3). As a result, a gender classifier trained on this data achieves at most 61.4% accuracy–the majority class (female) rate. Sim-PE iteratively refines selection from this noisy pool, ultimately choosing samples more aligned with target classes (Fig. 3), boosting accuracy by up to 21.1% (Tab. 1(a)).
Sim-PE maintains the strong data quality of PE. As shown in Tab. 1(b), Sim-PE and PE achieve similar FID. Unlike in MNIST (§ 4.2) where Sim-PE brought large gains, the modest improvement on CelebA stems from two factors. First, PE with foundation models already ranks 3rd in FID, leaving little room to improve. Second, Sim-PE here only selects from a fixed simulator-generated dataset. As seen in Fig. 3, these images differ from real CelebA (e.g., having larger faces), and Sim-PE cannot correct such discrepancies without access to simulator code. Having access to simulator code, as in § 4.2, could help alleviate such errors by allowing parameter adjustments. A hybrid approach combining foundation models and simulators, as we will explore next, may also offer further gains.
@@
(c) Sim-PE images ($\epsilon=10$)
Figure 4: The real and generated images on CelebA. The simulator is a weak rule-based avatar generator [19] significantly different from the real dataset. The top rows correspond to the “female” class, and the bottom rows correspond to the “male” class. The simulator generates images with incorrect classes. Sim-PE tends to generate faces with long hair for the female class and short hair for the male class (correctly), but the generated images have mode collapse issues.
@@ 4.4 Sim-PE with both Simulators and Foundation Models
In this section, we examine Sim-PE under weak simulators. As in the previous section, we use CelebA as the private dataset, but replace the simulator with a rule-based cartoon avatar generator [19]. As shown in Fig. 4, the generated avatars differ significantly from real CelebA images.
Sim-PE with weak simulators still learns useful features. From Tab. 2, we observe that downstream classifiers trained on Sim-PE with weak simulators achieve poor classification accuracy. However, two interesting results emerge: (1) Despite the significant difference between avatars and real face images, Sim-PE still captures certain characteristics of the two classes correctly. Specifically, Sim-PE tends to generate faces with long hair for the female class and short hair for the male class (Fig. 4). (2) Although the FID of Sim-PE is quite poor (Tab. 2), they still outperform many baselines (Tab. 1(b)). This can be explained by the fact that, as shown in [22], when DP noise is high, the training of many baseline methods becomes unstable. This results in images with noisy patterns, non-face images, or significant mode collapse, particularly for DP-NTK, DP-Kernel, and GS-WGAN. In contrast, Sim-PE is training-free, and thus it avoids these issues. See App. G for more result analysis.
Next, we explore the feasibility of using PE with both foundation models and the weak avatar simulator (§ 3.4). The results are shown in Tab. 2.
@@
Unlike other SoTA methods that are tied to a specific data synthesizer, this result suggests that PE is a promising framework that can easily combine the strengths of multiple types of data synthesizers.
@@
4.5 Efficiency of Sim-PE
Number of iterations. Like PE, Fig. 5 shows Sim-PE achieves good results in few iterations.
Computation cost. Since simulators could be much cheaper to run than foundation models, Sim-PE could be much more efficient than PE. For instance, on MNIST, each PE’s API call takes over 2400 GPU seconds, whereas Sim-PE takes less than 30 CPU seconds–an 80x speedup, not to mention the lower cost of CPU than GPU. See App. J for detailed results.
@@ 4.6 Validating the Design of Sim-PE
How does Sim-PE with simulator-generated data compare to other data selection algorithms? In § 3.3, we discussed two simple alternatives for simulator data selection. The comparison is in Tab. 3. We see that Sim-PE with iterative data selection outperforms the baselines on most metrics, validating the intuition outlined in § 3.3. However, the clustering approach used in the second baseline still has merit, as it results in a better FID for $\epsilon=10$. This idea is orthogonal to the design of Sim-PE and could potentially be combined for further improvement. We leave this exploration to future work.
@@
Due to space constraints, we defer the ablation studies on important parameters of Sim-PE and the impact of the alignment between the simulator and the private dataset to App. H.
@@ 5 Limitations and Future Work
In this paper, we demonstrate the potential of the PE framework for utilizing powerful simulators in DP image synthesis. While the results are promising, several important questions remain open:
@@
    Since the approach in § 3.3 can be applied to any public dataset (see App. F), Sim-PE could support applications such as pre-training data selection for private fine-tuning [67, 41].
@@
    This paper applies Sim-PE on images. In domains like networking and systems, simulators are more common than foundation models, and Sim-PE could offer even greater potential.
@@
    Sim-PE for image synthesis is still outperformed by the best baseline. Exploring better ways to combine simulators and foundation models could further push the limits of the PE framework.
  • $\bullet$
    It remains an open question how to effectively apply PE in domains where suitable foundation models, simulators, and public datasets are all unavailable.
Acknowledgement
The authors would like to thank the anonymous reviewers for their valuable feedback and suggestions.
This paper is the full version of our previous workshop papers [44, 45].
Appendix A The Prevalence and Importance of Simulators
In this paper, we define simulators as any data synthesizers that do not rely on neural networks. These simulators are widely used across various applications due to their unique advantages over neural network-based approaches.
The prevalence of simulators. Despite the widespread adoption of foundation models, simulators remain highly prevalent. Here are a few notable examples:
The importance of simulators. As demonstrated above, simulators continue to play a crucial role in industry applications. Even synthetic data generation also frequently relies on simulators rather than foundation models (e.g., Meta-Sim [38], DigiFace-1M [4]). This preference stems from several unique advantages of simulators over foundation models:
  • $\bullet$
    Rich annotations: Simulators provide additional structured labels due to explicit control over the data generation process. For instance, besides face images, Face Synthetics [61] offers pixel-level segmentation masks (e.g., identifying eyes, noses), which are highly valuable for downstream tasks such as face parsing.
  • $\bullet$
    Task-specific strengths: Certain generation tasks remain challenging for foundation models. For example, despite recent advances, foundation models still struggle with generating images containing text [54], whereas simulators can handle it easily. Our MNIST experiment was specifically designed to highlight this distinction.
  • $\bullet$
    Domain-specific strengths: In domains like networking, where robust foundation models are lacking, network simulators such as ns-3888https://www.nsnam.org/ remain a more reliable and scalable solution.
  • $\bullet$
    Greater reliability and control: Because simulators model the data generation process, they mitigate issues such as generating anatomically incorrect images (e.g., faces without noses [^9] https://www.reddit.com/r/StableDiffusion/comments/1e7dd62/weird_and_distorted_images_with_a1111_sd3/).
Appendix B Private Evolution
Alg. 1 presents the Private Evolution (PE) algorithm, reproduced from [46]. This algorithm represents the conditional version of PE, where each generated image is associated with a class label. It can be interpreted as running the unconditional version of PE separately for each class (Alg. 1).
@@ 1 $S_{\mathrm{syn}}\leftarrow\emptyset$
4       $S_{0}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$
@@
             $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples,S_{t-1},\sigma,H\right)$
@@
6             $S_{t}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
@@ 4       $histogram[i]\leftarrow histogram[i]+1$
Appendix C Justifying the Methodological Design of Sim-PE with Simulator-generated Data
In this section, we provide experimental evidence to support the claim in § 3.3: “If we already know that a sample $z_{i}$ is far from the private dataset, then its nearest neighbors in $S_{\text{sim}}$ are also likely to be far from the private dataset.”
Consider Sim-PE’s synthetic dataset $S_{t}$ from the $t$-th iteration. For each private sample $x_{i}$ in the private dataset $S_{\mathrm{priv}}{}$, we find its nearest sample in $S_{t}$ as $z_{i}$. $S(selected)_{t}=\left\{z_{i}\right\}_{i=1}^{N_{\mathrm{priv}}{}}$ contains samples close to $S_{\mathrm{priv}}{}$ (chosen by DP_NN_HISTOGRAM (Alg. 2) in PE under non-DP settings). Conversely, the remaining samples, $S(unselected)_{t}=S_{t}\setminus S(selected)_{t}$ are farther from $S_{\mathrm{priv}}{}$.
Now, we need to confirm that $S(unselected)_{t}$’s nearest neighbors in $S_{\text{sim}}{}$ are farther away from the private dataset $S_{\mathrm{priv}}{}$ than $S(selected)_{t}$. We define:
  • $\bullet$
    $z(Y)_{i}$ as the nearest neighbor of the private sample $x_{i}$ in $S(Y)_{t}$, for $Y\in\left\{selected,unselected\right\}$
  • $\bullet$
    $q(Y)_{j}^{i}$ as the $j-th$ nearest neighbor of $z(Y)_{i}\text{ in }S_{\text{sim}}$
  • $\bullet$
    $FID(Y)$ as the FID between $S_{priv}$ and $\left\{q(Y)_{j}^{i}\right\}_{i=1,j=1}^{N_{\mathrm{priv}},\gamma}$.
Fig. 6 shows that $FID(selected)$ is smaller (better) than $FID(unselected)$ across most PE iterations, supporting our claim.
Figure 6: FID of the nearest neighbors of the selected and unselected samples.
Appendix D Theoretical Analysis
Convergence analysis. Since Sim-PE only changes PE’s RANDOM_API and VARIATION_API, the original convergence analysis (App. E in [46]) remains valid for Sim-PE. Specifically:
  • $\bullet$
    The analysis assumes that RANDOM_API generates samples within a ball of diameter $D$ covering the private samples. Changing RANDOM_API affects $D$ but not the overall analysis procedure.
  • $\bullet$
    The analysis assumes that VARIATION_API draws samples from a Gaussian distribution—an essential assumption for any tractable analysis involving complex, unknown foundation models and simulators. Under this assumption, modifying VARIATION_API does not impact the analysis procedure.
Privacy analysis. Since Sim-PE does not change the way PE utilizes private data, the privacy analysis of PE [46] applies to Sim-PE.
Appendix E Discussion on the Chosen Metrics
In the main experiments, we use FID and the accuracy of downstream classifiers as the metrics.
Why we pick these metrics. We chose FID and classification accuracy because they are the most widely used metrics in DP image synthesis. Most of the baselines we compared to utilize only these two metrics or just one of them [22]. The only exception is GS-WGAN, which also uses Inception Score (IS) [55]. However, IS is more relevant for natural images (e.g., ImageNet images) because it evaluates the diversity across ImageNet classes, making it unsuitable for the datasets we considered in the main paper. Additionally, for the downstream task (ten-class classification for MNIST and binary classification for CelebA), we also follow prior works [22].
The differences between these metrics. FID measures the “fidelity” of synthetic data by mapping both synthetic and private data to an embedding space, approximating each with a Gaussian distribution, and calculating the Wasserstein distance between the two distributions. Classification accuracy measures the “utility” of synthetic data by evaluating its performance when used to train a downstream classifier, simulating real-world use where users expect the synthetic data to support good classifier performance on real data. FID and classification accuracy are complementary: FID focuses on distribution-level closeness, while classification accuracy is sensitive to outliers or samples near the classification boundary. A high score in one does not necessarily correlate with a high score in the other. This is why we often see cases where the best methods for these metrics differ [22].
Appendix F Sim-PE with CIFAR10 as the Private Dataset and ImageNet as the Simulator-generated Dataset
In this section, we apply the algorithm in § 3.3 on the setting where CIFAR10 is treated as the private dataset and ImageNet is treated as the simulator-generated dataset. We conduct this experiment for the following reasons:
  • $\bullet$
    In § 3.3, we discuss that Sim-PE can apply to any public dataset beyond simulator-generated data. This experiment provides evidence for that claim.
  • $\bullet$
    The original PE approach on CIFAR10 [46] uses a foundation model pretrained on ImageNet, which aligns with the simulator data in this experiment. Comparing their performance removes the impact of distribution alignment and highlights the differences in the RANDOM_API and VARIATION_API designs of PE and Sim-PE.
Experiment settings. We use $\epsilon=10$. The downstream task is classifying images into the 10 classes in CIFAR10, following DPImageBench [22] and the original PE paper [46].
Results. Sim-PE achieves FID=8.76, and acc=70.06%. In comparison, PE with a foundation model achieves similarly with FID=9.2 and acc=75.3%. These results suggest that Sim-PE and PE can be competitive when both suitable simulators and foundation models are available for the private dataset.
Appendix G More Analysis on Sim-PE with Weak Simulators
In Fig. 4, we see that Sim-PE’s generated images have many duplicates. In this section, we conduct more analysis on this phenomenon.
Fig. 7 shows that FID between generated images and the private dataset decreases (improves) across iterations, confirming that generated images better align with the private distribution over iterations. However, the number of unique samples in the generated dataset drops significantly (Fig. 8), leading to duplicates in Fig. 4. This may be due to the simulator being too weak, generating most images far from the private dataset. As a result, Sim-PE converges to a small set of good images.
Figure 7: Sim-PE (using DigiFace-1M)’s FID on CelebA improves over the course of the PE iterations.
Figure 8: The number of unique generated samples in each iteration of Sim-PE (using DigiFace-1M) on CelebA.
Appendix H Additional Ablation Studies
H.1 Sensitivity Analysis of $\alpha,\beta,\gamma$
These hyperparameters (defined in § 3.2 and 3.3) control the variation degree in VARIATION_API. Higher values induce larger variations. These are the only new hyperparameters introduced in Sim-PE.
Following the PE paper [46], we gradually decrease them across PE iterations. The idea is that larger variations at the beginning help Sim-PE/PE explore and find good seeds, while smaller variations are needed later for convergence. In our experiments, most are set to follow an arithmetic or geometric sequence across PE iterations (App. I) based on heuristics. In the following, we modify their values: for each hyperparameter, we either fix it to the largest or smallest value in its base sequence. Figs. 9 and 10 show that:
  • $\bullet$
    Setting these hyperparameters to a fixed large value is not ideal. For instance, setting a large text variation throughout greatly hurts FID and accuracy. Conversely, using a fixed small value has a smaller impact.
  • $\bullet$
    The default parameters do not always yield the best result (e.g., setting $\gamma$ to a fixed small value improves FID). This suggests that our main results in the paper could be improved with better hyperparameters.
(a) FID
(b) Classification accuracy
Figure 9: Accuracy and FID on MNIST with different $\alpha,\beta$ schedules ($\epsilon=10$). The “base” parameter schedule across PE iterations is: (a) font size’s $\alpha=[5,4,3,2]$, (b) rotation degree’s $\alpha=[9,7,5,3]$, (c) stroke width’s $\alpha=[1,1,0,0]$, (d) font’s $\beta=[0.8,0.4,0.2,0.0]$, (e) text’s $\beta=[0.8,0.4,0.2,0.0]$. “Small” (“large”) means fixing $\alpha/\beta$ to be the smallest (largest) value across PE iterations. For example, the unit under “font variation” and “large” indicates the results when we set font’s $\beta=[0.8,0.8,0.8,0.8]$.
(a) FID
(b) Classification accuracy
Figure 10: Accuracy and FID on CelebA using simulator-generated data with different $\gamma$ schedules ($\epsilon=10$). The “base” parameter schedule across PE iterations is $[1000,500,200,100,50,20]$. “Small” (“large”) means fixing $\gamma$ to be the smallest (largest) value across PE iterations. For example, the unit under “large” indicates the results when we set $\gamma=[1000,1000,1000,1000,1000,1000]$.
H.2 The Impact of Distribution Alignment
On CelebA, we show that Sim-PE with a simulator that aligns well with the private data outperforms standard PE (§ 4.3 and 1). However, Sim-PE with a simulator less aligned with the private data performs worse than PE (§ 4.4 and 2).
To provide a more controlled experiment, we fix the simulator type and only vary the alignment between the simulator and private data.
Experiment settings. We use the same 1.2M simulator-generated images from a computer graphics renderer as in § 4.3 and artificially adjust their alignment to CelebA. Specifically, for each image, we compute its distance to the closest CelebA image in the Inception embedding space, sort all images by this distance, and divide them into five subsets, $D_{0},...,D_{4}$ (e.g., the closest 0.24M images form $D_{0}$). Fig. 11 confirms that the FID between $D_{i}$ and CelebA increases with $i$, meaning that $D_{0}$ is the most aligned and $D_{4}$ the least. We then apply Sim-PE to each $D_{i}$ independently.
Results. Fig. 12 shows that as alignment decreases, the sample quality generally drops. Specifically, Sim-PE with $D_{0},...,D_{3}$ yields better classification accuracy than PE, while $D_{4}$ yields worse results than PE. This confirms that Sim-PE’s performance is influenced by the degree of alignment. Additionally, in all cases, Sim-PE’s FID is better than the original simulator data’s FID, demonstrating Sim-PE’s ability to select useful samples.
Figure 11: FID between the 5 constructed subsets of DigiFace-1M and CelebA.
(a) FID
(b) Classification accuracy
Figure 12: Results of Sim-PE-generated data using the subsets in Fig. 11 ($\epsilon=10$).
H.3 Class Label Information from the Simulators
For simulator 1, the target class label (i.e., the digit) is fully controlled by one parameter. For simulators 2 and 3, the target class label (i.e., the gender) is not directly controlled by any parameter, but could potentially be obtained by an external image gender classifier. One benefit of using domain-specific simulators is that we can potentially use the class label information to enhance data quality [61, 4]. To get a more comprehensive understanding of Sim-PE, we consider two settings: (1) Class label information is unavailable (abbreviated as “ClassUnavail”). We artificially make the problem more challenging by assuming that the class label information is not available. Therefore, Sim-PE has to learn to synthesize images with the correct class by itself. Our main experiments are based on this setting. (2) Class label information is available (abbreviated as “ClassAvail”). On MNIST, we further test how Sim-PE can be improved if the class label information is available. In this case, the RANDOM_API and VARIATION_API (Eqs. 1 and 2) are restricted to draw parameters from the corresponding class (i.e., the digit is set to the target class).
Results: Class label information from the simulators can be helpful. The results are presented in Tabs. 4 and 13. We observe that with digit information, the simulator-generated data achieves much higher classification accuracy (92.2%), although the FID remains low due to the generated digits exhibiting incorrect characteristics (Fig. 13). The fact that Sim-PE outperforms the simulator in both FID and classification accuracy across all settings suggests that Sim-PE effectively incorporates private data information to enhance both data fidelity and utility, even when compared to such a strong baseline. As expected, Sim-PE under ClassAvail matches or surpasses the results obtained in ClassUnavail across all settings, suggesting the usefulness of leveraging class label information.
(a) Real (private) images
(b) Simulator-generated images
(c) Sim-PE images ($\epsilon=10$)
Table 4: Accuracy (%) of classifiers trained on synthetic images and FID of synthetic images on MNIST under the “ClassAvail” setting. See Tab. 1 for results under the “ClassUnavail” setting for reference.
Algorithm FID $\downarrow$ Classification Acc. $\uparrow$
$\epsilon=1$ $\epsilon=10$ $\epsilon=1$ $\epsilon=10$
Simulator 86.0 ($\epsilon=0$) 92.2 ($\epsilon=0$)
Sim-PE 20.7 8.6 93.9 95.5
Appendix I Experimental Details